# Scaling Laws Revisited: Modeling the Role of Data Quality in Language Model Pretraining

### v1

@@ Abstract
Scaling laws for language model training traditionally characterize how performance scales with model size and dataset volume. Prior work has explored architecture variants and data treatments such as dataset filtering and noise injection in language model pretraining; however, these studies have not formalized data quality within a principled scaling law. We introduce a dimensionless data-quality parameter $Q$, and propose a quality-aware scaling law extending the Chinchilla framework to predict loss as a joint function of model size, data volume, and data quality. The law is motivated by an effective-sample-size and information-theoretic view of noisy or redundant corpora, and it admits two practical estimators for $Q$: (i) a corruption rate proxy and (ii) a deficiency measure. Through synthetic experiments in neural machine translation and autoregressive modeling—where we systematically control data quality via multiple levels of noise injection and coverage variation—we show that loss scales predictably with data quality and that higher-quality data can substantially reduce model size and hence compute requirements. Our results demonstrate a sublinear decay of effective data with quality and robustness to moderate data corruption; out-of-sample evaluations further validate the predictive form of the law. Unlike prior empirical analyses, our work establishes an explicit, generalizable law for data quality, offering concrete guidance for balancing data curation effort and model scale in large-scale pretraining.
@@ 1 Introduction
It is well understood in large-language model (LLM) training that both the amount of training data and its quality influence pretraining loss. Yet most existing scaling laws are formulated solely in terms of dataset size, remaining agnostic to quality [19, 16, 23, 35, 1, 8, 36, 41]. This leaves a gap between common intuition–“cleaner data trains better models”–and the quantitative laws that guide large-scale training.
The importance of such a quantitative framework is especially clear in specialized domains–business, scientific, or medical applications–where training corpora are limited in quantity and heterogeneous in quality. With the same compute budget, outcomes may differ drastically depending on how clean and representative the data are. There have been quite a number of studies to classify data quality among multiple “dimensions” [39, 14, 33, 22]. While the exact taxonomy varies, the core dimensions typically include: accuracy, completeness, consistency, timeliness, uniqueness, validity. The focus of these studies is usually datasets containing structured data that is clean and curated for a certain task or purpose. In contrast, when collecting data to pretrain large language models, large amounts of “found” data are collected and cleaned, usually by removing duplicate data and obviously “poor” quality data.
@@ 2 Related Works
Data quality has long been studied in information systems, where datasets are evaluated along multiple dimensions—accuracy, completeness, consistency, timeliness, uniqueness and validity—to ensure that structured records support downstream tasks. A consistent finding from this literature is that diverse imperfections degrade the utility of data and therefore motivate extensive curation practices [6]. Large-scale pretraining, however, relies on corpora assembled from the web. Here, quality varies widely across sources, and subtle forms of corruption or deficiency persist despite deduplication or heuristic filtering. Nonetheless, the key insight from information systems continues to hold: the quality of training data directly influences the performance of learned models.
@@
In the era of LLMs, quality has become a first-class design concern. Major corpora such as The Pile [15], RefinedWeb [32], and Dolma [34] highlight different philosophies of data construction, from assembling diverse sources to aggressively filtering and refining web text. Deduplication strategies [27] and contamination checks [30] have been shown to substantially affect evaluation reliability. Empirical analyses further suggest that careful filtering yields performance gains comparable to scaling compute, therefore implying the nontrivial role of quality in driving LLM performance.
@@
Alongside these developments, foundational studies on scaling laws formalized how performance scales with model size, dataset size, and compute. Kaplan et al. [25] showed that loss decreases predictably as models and datasets grow, while Hoffmann et al. [20] refined these results by introducing compute-optimal training. Despite their influence, these existing scaling laws generally assume that the underlying data is of fixed quality. In real-world corpora, noise, redundancy, or domain imbalance systematically distort scaling behavior, highlighting the need to explicitly account for quality.
@@
Several recent works have begun to address this gap. Bansal et al. [5] study scaling laws for neural machine translation, and show that while test loss scales predictably with dataset size, architectural choices and moderate noise primarily shift the scaling curves without changing the exponent. Goyal et al. [17] argue that data filtering is not compute-agnostic: the value of higher-quality data depends on the available compute budget, since reusing small clean datasets leads to diminishing returns. Li et al. [28] assess data quality by comparing perplexity gaps between models of different sizes, theoretically linking their metric to inverse scaling laws. These studies highlight the importance of data quality but stop short of providing a unified predictive law that couples quality with model size and dataset size. By introducing a dimensionless quality parameter $Q$ into scaling laws, our work formalizes the influence of data quality and offers predictive guidance for building smaller, high accuracy models in specialized domains.
Beyond these empirical analyses, several recent studies refine scaling laws from complementary perspectives. Chang et al. [9] propose parameter-constrained scaling laws that combine text diversity and syntheticity, revealing that effective training tokens depend on both data quantity and diversity. On the theoretical front, Bahri et al. [3] present a unifying theory of neural scaling laws, distinguishing variance-limited and resolution-limited regimes and connecting scaling exponents to generalization error. Although these contributions deepen our understanding of scaling, they do not incorporate an explicit scalar quality parameter. Our work complements this literature by modeling data quality via $Q$, linking it to effective sample size and generalization theory, and providing a unified scaling law that accounts for quality.
@@ 3.1 Data Corruption Rate
As an example, if we have a dataset consisting of $D$ tokens and $10\%$ are corrupted, then we would say that the $\text{CR}=10\%$ and the data quality $Q$ is $90\%$. To simplify our formulas, we assume that the data corruption rate is strictly less than $1$, so that $Q$ is strictly greater than 0. Note that standard sampling methods (e.g. [12]) can be used to estimate data corruption.
@@ 3.2 Data Deficiency Measures
We define the data deficiency data quality $Q(\omega)$ of a dataset $\omega$ by
@@ 4 Quality-Aware Scaling Laws
A widely used empirical scaling law for LLMs (also known as the Chinchilla Scaling Law by Hoffmann et al. [20]) is
@@
For example, one formulation studied in Hoffmann et al. [20] is:
@@ 4.1 Effective Sample Size
This assumption is consistent with classical results: in PAC learning with random classification noise of rate $\eta<1/2$, the sample complexity inflates by a factor $(1-2\eta)^{-2}$ [2, 26]; in regression with additive Gaussian noise of variance $\sigma^{2}$, effective samples scale with the signal-to-noise ratio $\mathrm{SNR}=1/\sigma^{2}$ [38]; and in information theory, channel capacity arguments show that mutual information is reduced by a multiplicative factor depending on the corruption level [10]. In the following two lemmas (proof deferred to the appendix), we show that the usable sample size can be written as $D_{\mathrm{eff}}=D\cdot g(Q)$ with $g(Q)$ well-approximated by a power law $Q^{\gamma}$ over practical ranges of $Q$:
@@ 4.2 Quality-Aware Scaling Law
with $\rho(1)=1$, $\rho(Q)$ monotone, and locally $\rho(Q)\approx c\,Q^{\gamma}$ as $Q\to 1$. Here $I(X;Z)$ denotes mutual information as usual. If the data-dependent loss scales as [^1] This assumption is motivated by information-theoretic generalization bounds [42, 7], which relate generalization error to the mutual information between training data and learned representations or weights, as well as by the Information Bottleneck principle [37].
@@ 5 Experimental Studies
To establish the quality aware scaling laws, we train decoder only language models across two tasks, neural machine translation (NMT) and causal language modeling (CLM). To capture the effects of quality and dataset size we train on 3 different dataset volumes and 7 degrees of quality. We use subsets of Paracrawl v8 ([4]) and C4 ([13]) as our training data for the translation and language tasks respectively. See Table 1. We optimize cross-entropy loss averaged over the relevant context and report the same on held out test datasets to measure the effectiveness of our quality aware scaling law for predicting the loss on out of sample data (Table 2. We use the parametric loss fitting approach described in [20] to find our parameters using two methods, Least Squares and Huber [21]. The full details of the experiments, and model training details can be found in the appendix.
@@ 5.1 Task Description
Our first task is English to German translation using the Paracrawl v8 English to German dataset [4]. We filter the raw data with min hash deduplication, language ID filtering with a 0.8 threshold and length filtering ($\leq$ 258) leaving $\sim$ 101M sentence pairs. We use a 8L GPT Neo model with a hidden size of 1024 with approximately $\sim$ 133M params. We also train our own BPE tokenizer with a vocabulary size of 32000. We train 3 replicates for each combination of the 3 datasets of size 500K, 1M and 2M sentence pairs and 7 quality levels for a total of 63 experiment runs. We train using AdamW with a maximum learning rate of 5e-4 and cosine decay with a 20% warmup ratio. The full model configuration and training hyper-parameters are available in the appendix.
@@
For the CLM objective we train our models on a subset of C4 (en) [13]. We run our pretraining experiments on a 8L Llama 3 [18] model with a hidden size of 512 and a context length of 2048. We do not use ROPE scaling. We employ random data truncation on the base dataset to sample within the context length. We also employ the pre-trained Llama-3.2-1B tokenizer. We sample datasets of three sizes; 100M , 1B and 10B tokens and train for a single epoch using fused AdamW [29] with a maximum learning rate of 1e-3 with cosine decay and a 10% warm-up ratio. We duplicate our experiment setting to match a total of 63 runs. The full model configuration and training hyperparameters are available in the appendix.
@@ 5.4 Experimental Setup
TASK METHOD B $\bm{\beta}$ $\bm{\gamma}$ $\bm{E^{+}}$
@@
TASK METHOD B $\bm{\beta}$ $\bm{\gamma}$ $\bm{E^{+}}$
@@ 5.5 Main Results
Our main results are summarized in Table 1, which estimates $B$, $\beta$, $\gamma$, and $E^{+}$, and Figure 1, which shows the test loss as $Q$ varies from $0.5$ to $1.0$. Predictably, we find that test loss $L$ decreases as we increase the volume of data $D$ and increase the fraction of high quality samples $Q$. Table 2 shows us that using our pre-trained models to evaluate loss on unseen data also follows a scaling law similar to our fit on in distribution data, thus showing evidence for scaling generalization. To further illustrate trade-offs between $D$ and $Q$, Figure 2 presents iso-loss contours, showing how data quality improvements can substitute for increased dataset size at fixed model capacity.
@@ 6 Summary and Conclusion
This paper revisits scaling laws for large language model pretraining by introducing an explicit and dimensionless measure of data quality, alongside traditional factors like model size and dataset volume. We propose a new quality-aware scaling law that extends the widely used Chinchilla framework, predicting pretraining loss as a joint function of model size, data volume, and data quality. By systematically injecting synthetic noise and varying coverage in neural machine translation and causal language modeling experiments, we demonstrate that higher data quality leads to significantly lower loss for a given model size and dataset. Notably, their findings show that for high-quality datasets, smaller models and less compute are needed to achieve strong results, which is highly relevant for domain-specific applications. The study provides a formal scaling law:
@@
where $Q$ measures data quality and $\gamma$ is empirically estimated. This extends earlier work that primarily focused on model and data size. Their experiments on both NMT and CLM tasks, using real datasets with systematically controlled levels of data corruption, confirm the predictive power of the new scaling law and offer concrete guidance for balancing data curation versus model scale. The work establishes a unified framework that quantifies data quality’s influence, enabling principled decisions for specialized LLM development in fields building smaller models over domain specific datasets, such as those that arise in business, scientific and medical applications.

### v2

@@ Abstract
Scaling laws for language model training traditionally characterize how performance scales with model size and dataset volume. Prior work has explored architecture variants and data treatments such as dataset filtering and noise injection in language model pretraining; however, these studies have not formalized data quality within a principled scaling law. We introduce a dimensionless data-quality parameter $Q$, and propose a quality-aware scaling law extending the Chinchilla framework to predict loss as a joint function of model size, data volume, and data quality. The law is motivated by an effective-sample-size and information-theoretic view of noisy or redundant corpora, and it admits two practical estimators for $Q$: (i) a corruption rate proxy and (ii) a deficiency measure. Through synthetic experiments in neural machine translation and autoregressive modeling—where we systematically control data quality via multiple levels of noise injection variation—we show that loss scales predictably with data quality and that higher-quality data can substantially reduce model size and hence compute requirements. Our results demonstrate a sublinear decay of effective data with quality and robustness to moderate data corruption; out-of-sample evaluations further validate the predictive form of the law. Unlike prior empirical analyses, our work establishes an explicit, generalizable law for data quality, offering concrete guidance for balancing data curation effort and model scale in large-scale pretraining.
@@ 1 Introduction
It is well understood in large-language model (LLM) training that both the amount of training data and its quality influence pretraining loss. Yet most existing scaling laws are formulated solely in terms of dataset size, remaining agnostic to quality (Henighan et al., 2020; Ghorbani et al., 2021; Ivgi et al., 2022; Sorscher et al., 2022; Alabdulmohsin et al., 2022; Caballero et al., 2023; Tao et al., 2024; Wu et al., 2025). This leaves a gap between common intuition–“cleaner data trains better models”–and the quantitative laws that guide large-scale training.
The importance of such a quantitative framework is especially clear in specialized domains–business, scientific, or medical applications–where training corpora are limited in quantity and heterogeneous in quality. With the same compute budget, outcomes may differ drastically depending on how clean and representative the data are. There have been quite a number of studies to classify data quality among multiple “dimensions” (Wang and Strong, 1996; Fox et al., 1994; Sidi et al., 2012; ISO/IEC, 2008). While the exact taxonomy varies, the core dimensions typically include: accuracy, completeness, consistency, timeliness, uniqueness, validity. The focus of these studies is usually datasets containing structured data that is clean and curated for a certain task or purpose. In contrast, when collecting data to pretrain large language models, large amounts of “found” data are collected and cleaned, usually by removing duplicate data and obviously “poor” quality data.
@@ 2 Related Works
Data quality has long been studied in information systems, where datasets are evaluated along multiple dimensions—accuracy, completeness, consistency, timeliness, uniqueness and validity—to ensure that structured records support downstream tasks. Note that what call completeness is also called coverage and closely to what is usually called diversity. A consistent finding from this literature is that diverse imperfections degrade the utility of data and therefore motivate extensive curation practices (Batini et al., 2009). Large-scale pretraining, however, relies on corpora assembled from the web. Here, quality varies widely across sources, and subtle forms of corruption or deficiency persist despite deduplication or heuristic filtering. Nonetheless, the key insight from information systems continues to hold: the quality of training data directly influences the performance of learned models.
@@
In the era of LLMs, quality has become a first-class design concern. Major corpora such as The Pile (Gao et al., 2020), RefinedWeb (Penedo et al., 2023), and Dolma (Soldaini et al., 2024) highlight different philosophies of data construction, from assembling diverse sources to aggressively filtering and refining web text. Deduplication strategies (Lee et al., 2022) and contamination checks (Magar and Schwartz, 2022) have been shown to substantially affect evaluation reliability. Empirical analyses further suggest that careful filtering yields performance gains comparable to scaling compute, therefore implying the nontrivial role of quality in driving LLM performance.
@@
Alongside these developments, foundational studies on scaling laws formalized how performance scales with model size, dataset size, and compute. Kaplan et al. (2020) showed that loss decreases predictably as models and datasets grow, while Hoffmann et al. (2022) refined these results by introducing compute-optimal training. Despite their influence, these existing scaling laws generally assume that the underlying data is of fixed quality. In real-world corpora, noise, redundancy, or domain imbalance systematically distort scaling behavior, highlighting the need to explicitly account for quality.
@@
Several recent works have begun to address this gap. Bansal et al. (2022) study scaling laws for neural machine translation, and show that while test loss scales predictably with dataset size, architectural choices and moderate noise primarily shift the scaling curves without changing the exponent. Goyal et al. (2024) argue that data filtering is not compute-agnostic: the value of higher-quality data depends on the available compute budget, since reusing small clean datasets leads to diminishing returns. Li et al. (2024) assess data quality by comparing perplexity gaps between models of different sizes, theoretically linking their metric to inverse scaling laws. These studies highlight the importance of data quality but stop short of providing a unified predictive law that couples quality with model size and dataset size. By introducing a dimensionless quality parameter $Q$ into scaling laws, our work formalizes the influence of data quality and offers predictive guidance for building smaller, high accuracy models in specialized domains.
Beyond these empirical analyses, several recent studies refine scaling laws from complementary perspectives. Chang et al. (2024) propose parameter-constrained scaling laws that combine text diversity and syntheticity, revealing that effective training tokens depend on both data quantity and diversity. On the theoretical front, Bahri et al. (2024) present a unifying theory of neural scaling laws, distinguishing variance-limited and resolution-limited regimes and connecting scaling exponents to generalization error. Although these contributions deepen our understanding of scaling, they do not incorporate an explicit scalar quality parameter. Our work complements this literature by modeling data quality via $Q$, linking it to effective sample size and generalization theory, and providing a unified scaling law that accounts for quality.
@@ 3.1 Data Corruption Rate
As an example, if we have a dataset consisting of $D$ tokens and $10\%$ are corrupted, then we would say that the $\text{CR}=10\%$ and the data quality $Q$ is $90\%$. To simplify our formulas, we assume that the data corruption rate is strictly less than $1$, so that $Q$ is strictly greater than 0. Note that standard sampling methods (e.g. Deming (1966)) can be used to estimate data corruption.
@@ 3.2 Data Deficiency Measures
We define the data quality induced by data deficiency $Q(\omega)$ of a dataset $\omega$ by
@@ 4 Quality-Aware Scaling Laws
A widely used empirical scaling law for LLMs (also known as the Chinchilla Scaling Law by Hoffmann et al. (2022)) is
@@
For example, one formulation studied in Hoffmann et al. (2022) is:
@@ 4.1 Effective Sample Size
This assumption is consistent with classical results: in PAC learning with random classification noise of rate $\eta<1/2$, the sample complexity inflates by a factor $(1-2\eta)^{-2}$ (Angluin and Laird, 1988; Kearns, 1998); in regression with additive Gaussian noise of variance $\sigma^{2}$, effective samples scale with the signal-to-noise ratio $\mathrm{SNR}=1/\sigma^{2}$ (Tsybakov, 2009); and in information theory, channel capacity arguments show that mutual information is reduced by a multiplicative factor depending on the corruption level (Cover and Thomas, 2006). In the following two lemmas (proof deferred to the appendix), we show that the usable sample size can be written as $D_{\mathrm{eff}}=D\cdot g(Q)$ with $g(Q)$ well-approximated by a power law $Q^{\gamma}$ over practical ranges of $Q$:
@@ 4.2 Quality-Aware Scaling Law
with $\rho(1)=1$, $\rho(Q)$ monotone, and locally $\rho(Q)\approx c\,Q^{\gamma}$ as $Q\to 1$. Here $I(X;Z)$ denotes mutual information as usual. If the data-dependent loss scales as [^1] This assumption is motivated by information-theoretic generalization bounds (Xu and Raginsky, 2017; Bu et al., 2020), which relate generalization error to the mutual information between training data and learned representations or weights, as well as by the Information Bottleneck principle (Tishby et al., 1999).
@@
4.3 Comparison with Other Quality-aware Scaling Laws
Table 1: Relating different realizations of the data deficiency term $\Delta(\omega)$ to the literature
Definition of $\Delta(\omega)$ Interpretation of Data Deficiency Paper
$\mu_{1}E+\mu_{2}\frac{1}{F}+\mu_{3}G+\mu_{4}H$ (equation 3) Decomposition into noise ($E$), inverse coverage ($1/F$), and repeated-data measure ($G$), and syntheticity ($S$). Ours
$\mu_{2}\,\mathrm{Dis}(\rho)$ Cluster-based density $\rho$, with $\mathrm{Dis}(\rho)=-\ln(1-\rho)$. Small $\rho\!\to\!$ high-quality; large $\rho\!\to\!$ low-quality. Chen et al. (2025)
$\mu_{2}\,\mathrm{DR}(\omega)+\mu_{4}\,S(\omega)$ Compression-based density (compressibility as proxy for diversity) $\mathrm{DR}(\omega)=\mathrm{CR}^{-1}(\omega)$ and $\mathrm{CR}=\frac{\mathrm{size}}{\mathrm{compressed\ size}}$, and data syntheticity $S(\omega)$. Higher DR, higher $S$ $\to\!$ lower quality. Chang et al. (2024)
$\mu_{3}\frac{k}{\tau}$ Declining marginal value of data with repeated exposure over $k$ epochs; linear–exponential saturation depending on $\tau$. Models over-training on repeated data. Goyal et al. (2024)
To relate our data dispersion to the literature, assume that the data deficiency is of the form
$\Delta(\omega)=\mu_{1}E+\mu_{2}\frac{1}{F}+\mu_{3}G+\mu_{4}H,$ (3)
where $E$ quantifies the noise in the data, $F$ quantifies the coverage/diversity of the data, $G$ quantifies the amount of repeated data, and $H$ quantifies the loss in data quality due to synthetic data. As shown in Table 1, by assuming particular forms for each of these terms, we recover the data quality scaling laws in Chen et al. (2025), Goyal et al. (2024) and Chang et al. (2024).
@@ 5 Experimental Studies
To establish the quality aware scaling laws, we train decoder only language models across two tasks, neural machine translation (NMT) and causal language modeling (CLM). To capture the effects of quality and dataset size we train on 3 different dataset volumes and 7 degrees of quality. We use subsets of Paracrawl v8 (Bañón et al. (2020)) and C4 (Dodge et al. (2021)) as our training data for the translation and language tasks respectively. See Table 2. We optimize cross-entropy loss averaged over the relevant context and report the same on held out test datasets to measure the effectiveness of our quality aware scaling law for predicting the loss on out of sample data (Table 3. We use the parametric loss fitting approach described in Hoffmann et al. (2022) to find our parameters using two methods, Least Squares and Huber (Huber, 1992). The full details of the experiments, and model training details can be found in the appendix.
@@ 5.1 Task Description
Our first task is English to German translation using the Paracrawl v8 English to German dataset (Bañón et al., 2020). We filter the raw data with min hash deduplication, language ID filtering with a 0.8 threshold and length filtering ($\leq$ 258) leaving $\sim$ 101M sentence pairs. We use a 8L GPT Neo model with a hidden size of 1024 with approximately $\sim$ 133M params. We also train our own BPE tokenizer with a vocabulary size of 32000. We train 3 replicates for each combination of the 3 datasets of size 500K, 1M and 2M sentence pairs and 7 quality levels for a total of 63 experiment runs. We train using AdamW with a maximum learning rate of 5e-4 and cosine decay with a 20% warmup ratio. The full model configuration and training hyper-parameters are available in the appendix.
@@
For the CLM objective we train our models on a subset of C4 (en) (Dodge et al., 2021). We run our pretraining experiments on a 8L Llama 3 (Grattafiori et al., 2024) model with a hidden size of 512 and a context length of 2048. We do not use ROPE scaling. We employ random data truncation on the base dataset to sample within the context length. We also employ the pre-trained Llama-3.2-1B tokenizer. We sample datasets of three sizes; 100M , 1B and 10B tokens and train for a single epoch using fused AdamW (Loshchilov and Hutter, 2019) with a maximum learning rate of 1e-3 with cosine decay and a 10% warm-up ratio. We duplicate our experiment setting to match a total of 63 runs. The full model configuration and training hyperparameters are available in the appendix.
@@ 5.4 Experimental Setup
TASK METHOD B $\bm{\beta}$ $\bm{\gamma}$ $\bm{E}$
@@
TASK METHOD B $\bm{\beta}$ $\bm{\gamma}$ $\bm{E}$
@@ 5.5 Main Results
Our main results are summarized in Table 2, which estimates $B$, $\beta$, $\gamma$, and $E$, and Figure 1, which shows the test loss as $Q$ varies from $0.5$ to $1.0$. Predictably, we find that test loss $L$ decreases as we increase the volume of data $D$ and increase the fraction of high quality samples $Q$. Table 3 shows us that using our pre-trained models to evaluate loss on unseen data also follows a scaling law similar to our fit on in distribution data, thus showing evidence for scaling generalization. To further illustrate trade-offs between $D$ and $Q$, Figure 2 presents iso-loss contours, showing how data quality improvements can substitute for increased dataset size at fixed model capacity.
@@
Robustness and generalization.
We have included additional validations to verify the robustness of our findings (see Appendix E for full details). First, we confirmed that our synthetic noise model serves as a valid proxy for semantic degradation by observing a strictly monotonic decrease in embedding similarity with increasing noise (Figure 4). We performed learning rate sweeps and found that $\gamma$ is stable across practical learning rate regimes (Table 9). We provide a practical guide for estimating $Q$ in Section E.3.
@@ 6 Summary and Conclusion
This paper revisits scaling laws for large language model pretraining by introducing an explicit and dimensionless measure of data quality, alongside traditional factors like model size and dataset volume. We propose a new quality-aware scaling law that extends the widely used Chinchilla framework, predicting pretraining loss as a joint function of model size, data volume, and data quality. By systematically injecting synthetic noise in neural machine translation and causal language modeling experiments, we demonstrate that higher data quality leads to significantly lower loss for a given model size and dataset. Notably, our findings show that for high-quality datasets, smaller models and less compute are needed to achieve strong results, which is highly relevant for domain-specific applications. The study provides a formal scaling law:
@@
where $Q$ measures data quality and $\gamma$ is empirically estimated. This extends earlier work that primarily focused on model and data size. Our experiments on both NMT and CLM tasks, using real datasets with systematically controlled levels of data corruption, confirm the predictive power of the new scaling law and offer concrete guidance for balancing data curation versus model scale. The work establishes a unified framework that quantifies data quality’s influence, enabling principled decisions for specialized LLM development in fields building smaller models over domain specific datasets, such as those that arise in business, scientific and medical applications.
@@
Acknowledgments
This research was supported, in part, funded by the National Heart, Lung, and Blood Institute (NHLBI) of the National Institutes of Health (NIH). The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the NIH/NHLBI. We also acknowledge the University of Chicago’s Research Computing Center for their support of this work.